OpenStudy (anonymous):

ALGEBRA 2 ----------------------------------------***WILL BECOME FAN AND GIVE MEDAL*** Solve the following system of equations and show all work. y = −x2 + 4 y = 2x + 1

OpenStudy (plasmataco):

-x squared?

OpenStudy (anonymous):

yeah

OpenStudy (plasmataco):

oh.

OpenStudy (plasmataco):

substitute.

OpenStudy (anonymous):

just need the steps

OpenStudy (plasmataco):

replace the bottem equations with Y on the top equation.

OpenStudy (plasmataco):

add X^2 and subtract 4 from both sides to get X^2+2x-3

OpenStudy (anonymous):

alright got that

OpenStudy (plasmataco):

the simplify to get (X-1)(X+3)

OpenStudy (plasmataco):

X can be either 1 or -3.

OpenStudy (plasmataco):

now, sub it in.

OpenStudy (anonymous):

x^2 + 3x -x -3

OpenStudy (plasmataco):

yup

OpenStudy (plasmataco):

you got it.

OpenStudy (anonymous):

so that's the answer?

OpenStudy (plasmataco):

no.

OpenStudy (plasmataco):

simplify it to (X-1)(X+3)

OpenStudy (plasmataco):

then, for (X-1)(X+3)=0, X could be 1 or negative 3.

OpenStudy (anonymous):

x^2 + 3x -x - 3 x^2 + 2x -3

OpenStudy (plasmataco):

sub it in so Y=2X+1, X=1, Y=3. (1,3) when X=-3 Y=5 so (-3,5)

OpenStudy (plasmataco):

nononono

OpenStudy (plasmataco):

don't multiply both of them.

OpenStudy (plasmataco):

make separate equations like X-1=0 and X+3=0

OpenStudy (plasmataco):

then, look up to see the rest.

OpenStudy (plasmataco):

I'm 13 XD

OpenStudy (anonymous):

nice

OpenStudy (plasmataco):

well, get it?